Describe in words the graph of each of these curves below. Include in your description the shape, along with other possible relevant information such as length, width, and center points.
a. (X+2)2 + (Y-4)2 = 36
b. Y = 3X2
c. Y = X2 + X
d. (X-1)2 + (Y-8)2 = 16

Answer by stanbon(75887)

Describe in words the graph of each of these curves below. Include in your description the shape, along with other possible relevant information such as length, width, and center points.
a. (X+2)^2 + (Y-4)^2 = 36
circle; center at (-2,4); radius = 6
--------------------------------------
b. Y = 3X^2
parabola; vertex (0,0); opening up
--------------------------------------
c. Y = X^2 + X
parabola: vertex (-1/2); open up
------------------------------------
d. (X-1)^2 + (Y-8)^2 = 16
circle; center (1,8); radius = 4
==================================
Cheers,
Stan H.